OpenStudy (anonymous):

In 1990, Tom Jager of the United States set a world record by swimming 50M at a rate of 2.29 m/s. Previously, the fastest swimming rate on record was 2.26 m/s, set in 1975 by David Holmes Edgar, also of the United States. Calculate the percentage by which the rate increased

OpenStudy (immanuelv):

2.29-2.26/2.26=.0132743363---so the rate increased by 1.32743363%..

OpenStudy (anonymous):

hey Immanuelv

OpenStudy (immanuelv):

hi

OpenStudy (anonymous):

i kind got something different what did i do wrong would you mind explaining to me

OpenStudy (anonymous):

2.26/2.29=.986 .986(100)=98%

OpenStudy (immanuelv):

you subtract old record by the new one.. then divide by the new record first do the subtraction... then multiply the answer by 100 to get the percent!

OpenStudy (anonymous):

ok

OpenStudy (immanuelv):

did u get it right this time?

OpenStudy (anonymous):

ok i took 2.29/2.26-2.26=1.27

OpenStudy (immanuelv):

no.... its (2.29-2.26)/2.26

OpenStudy (anonymous):

ok i get it now i get o.0132743362 i them move the decennial making it 1.327....

OpenStudy (immanuelv):

no ... multiply by 100 to make it a percent.... hope it helped! :)

OpenStudy (anonymous):

oky thats where i missed up thks i get it now

OpenStudy (immanuelv):

no prob :)

OpenStudy (anonymous):

can i shoot you another one

OpenStudy (immanuelv):

i'll try to help!

OpenStudy (anonymous):

tks

OpenStudy (anonymous):

Use an inequality and the five step process to solve each problem The women’s volleyball team can spend at most $700.00 for its awards banquet at a local restaurant. If the restaurant charges a $100.00 set up-fee plus $24 per person, at most how many can attend?

OpenStudy (anonymous):

you know what dora told nemo's dad?

OpenStudy (anonymous):

dory